# Spatial Re-parameterization for N:M Sparsity

### v2

@@ 4.4 Performance Analysis
Additional Overhead. We further elucidate the overhead implicated in Floating Point Operations (FLOPs) and parameters concomitant with the introduction of SpRe. Table IX enumerates the quantitative outcomes from employing SR-STE for the sparsification of ResNet-50 on ImageNet-1K. The deployment of SpRe does, indeed, incur a modicum of additional overhead during the training phase, albeit minimal. However, during the inference (test) stage, the quantity of parameters and FLOPs perseveres consistently, enabling efficacious deployment and acceleration founded on the Sparse Tensor Core [12]. Concurrently, juxtaposed with the results presented in Table IX, SpRe achieves a palpable enhancement in performance with the introduction of merely a negligible training overhead, affirming its undoubted practical utility.
TABLE IX: FLOPs and parameters comparison for pruning ResNet-50 on ImageNet-1K at training and testing stages.
@@ 5 Limitation
We further discuss unexplored limitations, which will be our future focus. SpRe is particularly presented for N:M sparsity in convolutional neural networks upon our observations of spatial sparsity variability. Although not the focus of this current work, it would be interesting for future work to examine similar discoveries to drive further enhancement for N:M sparsity in networks with different typologies, e.g., Vision Transformers (ViTs). Besides, Though improving the performance, SpRe also brings some extra burdens for N:M sparsity in the training time. A promising direction in the future is to develop more efficient ways to mitigate the lack in spatial sparsity variability of N:M sparsity.
@@ Acknowledgement
This work was supported by National Key R&D Program of China (No.2022ZD0118202), the National Science Fund for Distinguished Young Scholars (No.62025603), the National Natural Science Foundation of China (No. U21B2037, No. U22B2051, No. 62176222, No. 62176223, No. 62176226, No. 62072386, No. 62072387, No. 62072389, No. 62002305 and No. 62272401), and the Natural Science Foundation of Fujian Province of China (No.2021J01002, No.2022J06001).

### v3

@@ 4.4 Performance Analysis
TABLE IX: Theoretical FLOPs and parameters comparison for pruning ResNet-50 on ImageNet-1K at training and testing stages.
@@
TABLE X: Training latency (A100 GPU hour/epoch) and Memory cost (GB) for pruning ResNet-50.
Method N:M Training Memory
Latency Cost
ResNet-50 - 0.39 27.9G
ASP 2:4 0.40 28.0G
w. SpRe 2:4 0.42 28.3G
SR-STE 2:4 0.41 28.2G
w. SpRe 2:4 0.44 28.6G
Additional Overhead. We further elucidate the overhead implicated in Floating Point Operations (FLOPs) and parameters concomitant with the introduction of SpRe. Table IX enumerates the theoretical results from employing SR-STE for the sparsification of ResNet-50 on ImageNet-1K at different N:M patterns. In addition, Table X report the training memory and speed results from employing ASP and SR-STE at 2:4 sparsity. The deployment of SpRe does, indeed, incur a modicum of additional overhead during the training phase, albeit minimal. However, during the inference (test) stage, the quantity of parameters and FLOPs perseveres consistently, enabling efficacious deployment and acceleration founded on the Sparse Tensor Core [12]. Concurrently, juxtaposed with the results presented in Table IX, SpRe achieves a palpable enhancement in performance with the introduction of merely a negligible training overhead, affirming its undoubted practical utility.
@@ 5 Limitation
We further discuss unexplored limitations, which will be our future focus. SpRe is particularly presented for N:M sparsity in convolutional neural networks upon our observations of spatial sparsity variability. Although not the focus of this current work, it would be interesting for future work to examine similar discoveries to drive further enhancement for N:M sparsity in networks with different typologies, e.g., Vision Transformers (ViTs), BERT. While non-CNN architectures do not inherently incorporate spatial sparsity, investigating the distributional disparities between unstructured and N:M sparsity patterns, and further designing corresponding re-parameterization mechanisms presents a promising approach to improve the performance of N:M sparse networks. Besides, though improving the performance, SpRe also brings some extra burdens for N:M sparsity in the training time. A promising direction in the future is to develop more efficient ways to mitigate the lack in spatial sparsity variability of N:M sparsity.
@@ Acknowledgement
This work was supported by National Key R&D Program of China (No.2022ZD0118202), the National Science Fund for Distinguished Young Scholars (No.62025603), the National Natural Science Foundation of China (No.624B2119, No. U21B2037, No. U22B2051, No. 62176222, No. 62176223, No. 62176226, No. 62072386, No. 62072387, No. 62072389, No. 62002305 and No. 62272401), and the Natural Science Foundation of Fujian Province of China (No.2021J01002, No.2022J06001).